OpenStudy (anonymous):

Use the technique developed in this section to solve the minimization problem. Minimize C = -2x + y Subject to x + 2y ( or _) 0, y (>or_) 0

OpenStudy (anonymous):

graph the 4 inequalities. They you will get a polygon. Find the vertices (corners). Substitute the x and y values of each into the C= the highest value and lowest value will be the max and min

OpenStudy (anonymous):

Not coming up right

OpenStudy (anonymous):

well you equation is missing 3x + 2y<=???

OpenStudy (anonymous):

12 is the missing number

OpenStudy (anonymous):

Here is the graph, only quadrant 1 the vertices are (0,3), (3,1.5), (4,0) and (0,0)

OpenStudy (anonymous):

c=-2x+y 0,3 = 3 maximum 3,1.5 = -4.5 4,0 = -8 minimum 0,0 = 0

OpenStudy (anonymous):

whenever I try to graph that one, it won't work properly.

OpenStudy (anonymous):

what does your graph look like. The equations are in standard form so use intercepts to graph x+2y=6 (6,0),(0,3) 3x+2y=12 (4,0),(0,6)

OpenStudy (anonymous):

Still so very lost...

OpenStudy (anonymous):

if the equation is in standard form ax + by = c the x-intercept is (c/a,0) and the y-intercept is (0,c/b) it makes it easier to graph using those two points

OpenStudy (anonymous):

Okay, i'm still trying..

OpenStudy (anonymous):

Oh well, going to have to fail this one